OpenStudy (callisto):

Polynomials #1 Question 41 It is given that f(3x) = 54x^3 -27x^2 +px +q. When f(x) is divided by (x-3), the remainder is 42. Find the remainder when f(x/3) is divided by x-9 *Note: I'm helping my sister but I'm in trouble too :|*

OpenStudy (anonymous):

I thought someone was typing an answer.

OpenStudy (cwtan):

TT bluring....... btw i found -4.... I duno whether it is right or wrong...

OpenStudy (callisto):

The answer is not -4 :|

OpenStudy (cwtan):

younger sister? it's complicated... LOL

OpenStudy (callisto):

Younger sister...

OpenStudy (anonymous):

p+q=6?

OpenStudy (callisto):

Nope :| p+q=15

OpenStudy (cwtan):

Big trouble

OpenStudy (anonymous):

oh yes 15

OpenStudy (anonymous):

i know wait

OpenStudy (anonymous):

:) why my answer is 42

OpenStudy (callisto):

It is!!!!

OpenStudy (anonymous):

ohhohohoh

OpenStudy (anonymous):

Lol this question is funny.

OpenStudy (callisto):

How did you get that answer? @Ackhat

OpenStudy (ash2326):

We have \[f(3x)=54x^3-27x^2+px+q\] It can be written as \[f(3x)=2(3x)^3-3(3x)^2+\frac p 3 (3x) +q\] so \[f(x)=2x^3-3x^2+\frac p 3 x+q\] We are given that when f(x) is divided by (x-3) the remainder is 42 so Using remainder theorem \[f(3)=42\] Now you can find P+q from here, Next find \(f(\frac x 3)\) to find the remainder when f(x/3) is divided by (x-9) put x=9 in f(x/3)

OpenStudy (anonymous):

f(x)=2x^3-3x^2+p/3x+q

OpenStudy (callisto):

.... I've thought about it.....I swear......

OpenStudy (anonymous):

f(x)/(x-3)=k+42/(x-3) 3f(x/3)/(x-9)= k + 3*42/(x-9)

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

i simpy divide by x-3 and got p+q=15 and after by x-9 and got that the reminder is p+q+27

OpenStudy (callisto):

Hold on... @Ishaan94 How does it work: 3f(x/3)/(x-9)= k + 3*42/(x-9)? @Ackhat Do you mean you divided the equation f(x) by (x-3) first, then divide f(x) by (x-9)? or ...?

OpenStudy (anonymous):

f(x)/(x-3) f(x/3)/(x-9)

OpenStudy (anonymous):

\[\frac{f(x)}{x-3} = P(x) + \frac{42}{x-3}\]\[x = \frac{x}{3}\] \[\frac{3\cdot f\left(x/3\right)}{x-9} = P(x) + \frac{42\cdot 3}{x-9}\] I love my solution <3

OpenStudy (callisto):

Okay, got it. My calculation mistake :| @Ackhat

OpenStudy (callisto):

@Ishaan94 More explanation is appreciated :) (sorry... I'm stupid :| )

OpenStudy (anonymous):

Beautiful solution Ishaan.

OpenStudy (anonymous):

Calli: Division Algorithm

OpenStudy (anonymous):

I am not a good teacher, Sorry. What part you didn't understand callisto?

OpenStudy (anonymous):

Thank you very much foolformath

OpenStudy (callisto):

\[\frac{3\cdot f\left(x/3\right)}{x-9} = P(x) + \frac{42\cdot 3}{x-9}\] ^ don't know where it comes..

OpenStudy (anonymous):

Bezu

OpenStudy (anonymous):

:)

OpenStudy (cwtan):

What a simple solution!!! Great job @Ishaan94 !!!!!

OpenStudy (anonymous):

Substitute \(x=\frac x3 \) in \[ \frac{f(x)}{x-3} = P(x) + \frac{42}{x-3} \]

OpenStudy (callisto):

Oh... Got it!!!! Thanks!!! Lovely solution :)

OpenStudy (anonymous):

Okay. \[\large\frac{f\left(\frac{x}{3}\right)}{\frac{x}3-3} = P + \frac{42}{\frac x3-3}\]Where P is any quadratic polynomial. \[\large \implies \frac{f\left(\frac x3\right)}{\frac{x-9}3} = P + \frac{42}{\frac {x -9}3}\]

OpenStudy (anonymous):

No, it is beautiful :|

OpenStudy (anonymous):

What's 'Bezu'? @Akchat

OpenStudy (anonymous):

@Ackhat *

OpenStudy (anonymous):

never mind

OpenStudy (anonymous):

ok this is a theorem

OpenStudy (callisto):

The most wonderful thing is that my sister understands it :) Once again, thank you everyone :)